OpenStudy (nerdsarecool):

A bookstore had 81 copies of a magazine. Yesterday it sold 7/9 of them. How many copies were sold yesterday?

OpenStudy (nerdsarecool):

Can someone help me with this

OpenStudy (nerdsarecool):

shishi990

OpenStudy (nerdsarecool):

shishi990

OpenStudy (anonymous):

7/9 of 81 is \[\frac{ 7 }{ 9 }\times 81\]

OpenStudy (prizzyjade):

\[81 \times \frac{ 7 }{ 9 }= 63\]

OpenStudy (mathstudent55):

\(\dfrac{7}{9} \times 81 = \dfrac{7}{9} \times \dfrac{81}{1} = \dfrac{7 \times 81}{9 \times 1} = \dfrac{7 \times \cancel{9} \times 9}{\cancel {9} \times 1} = 7 \times 9 = 63\)